Pratte Boat Wash's cost formula for its cleaning equipment and supplies is $2,550 per month plus $46 per boot. For the month of April, the company planned for activity of 60 boats, but the actual level of activity was 15 boats. The actual cleaning equipment and supplies for the month was $3,370. 

The activity variance for cleaning equipment and supplies in April would be closest to: 

Multiple Choice 

  • $1940 F 

  • $1940 U 

  • $2,070 U 

  • $2.070 F


Hirons Air uses two measures of activity, flights and passengers, in the cost formulas in its budgets and performance reports. The cost formula for plane operating costs is $59,450 per month plus $3,182 per flight plus $19 per passenger. The company expected its activity in November to be 82 flights and 252 passengers, but the actual activity was 85 flights and 254 passengers. The actual cost for plane operating costs in November was $323,610. The spending variance for plane operating costs in November would be closest to: 

Multiple Choice 

  • $1,552 F 

  • $11,136 F 

  • $11136 U

  • $1,5520 U


Petrus Framing's cost formula for its supplies cost is $1,760 per month plus $10 per frame. For the month of March, the company planned for activity of 616 frames, but the actual level of activity was 624 frames. The actual supplies cost for the month was $8,420. The activity variance for supplies cost in March would be closest to: 

Multiple Choice 

  • $500 F 

  • $500 U 

  • $80 F 

  • $80 U
Answer #1

1) Activity variance for cleaning equipment = Budgeted cost-Flexible budget cost = (60-15)*46 = 2070 F

So answer is d) $2070 F

2) Spending variance for plane operating cost = (59450+3182*85+19*254)-323610 = 11136 F

So answer is b) $11136 F

3) Activity variance = (616-624)*10 = 80 U

So answer is d) $80 U